solver for c: a=square root 2 over bc
Answer by edjones(8007)

a=sqrt(2)/bc
a(c/a)=(c/a)sqrt(2)/bc
c=(c/a)sqrt(2)*1/bc
c=sqrt(2)/ab
.
Ed